# https://m-niemeyer.github.io/radsplat/ RadSplat: Radiance Field-Informed Gaussian Splatting for Robust Real-Time Rendering with 900+ FPS

### v1

@@ 1 Introduction
While neural field methods have achieved unprecedented view synthesis quality even for challenging real-world captures [2, 32, 28], most approaches are limited by the high compute costs of volumetric rendering. In order to achieve real-time frame rates, recent works reduce network complexity [12, 35], cache intermediate outputs [56, 10], or extract 3D meshes [72, 41, 73, 53]. Nevertheless, all methods essentially trade reduced quality and increased storage costs for faster rendering, and are incapable of maintaining state-of-the-art quality in real-time – the goal of this work.
Recently, rasterization-based 3D Gaussian Splatting (3DGS) [21] has emerged as a natural alternative to neural fields. The representation admits real-time frame rates with view synthesis quality rivaling the state-of-the-art in neural fields. 3DGS, however, suffers from a challenging optimization landscape and an unbounded model size. The number of Gaussian primitives is not known a priori, and carefully-tuned merging, splitting, and pruning heuristics are required to achieve satisfactory results. The brittleness of these heuristics become particularly evident in large scenes where phenomena such as exposure variation, motion blur, and moving objects are unavoidable. An increasing number of primitives further leads to a potentially-unmanageable memory footprint and reduced rendering speed, strongly limiting model quality for larger scenes.
@@
Our method exhibits state-of-the-art reconstruction quality on both medium and large scenes, with PSNR up to 1.87 dB higher than 3DGS and SSIM exceeding Zip-NeRF, the current state-of-the-art in offline view synthesis (see Fig. 1). At the same time, our method renders up to 907 frames per second, over 3.6 faster than 3DGS and more than 3,000 faster than Zip-NeRF.
@@ 2 Related Work
Neural Fields. Since their introduction in the context of 3D reconstruction [31, 42, 6, 69], neural fields have become one of the most promising methods for many 3D vision tasks including 3D/4D reconstruction [73, 25, 44, 39, 43, 46], 3D generative modeling [59, 38, 47, 4, 27, 64], and view synthesis [32, 1, 2]. Key to their success is among others simplicity, state-of-the-art performance, and robust optimization [69]. In contrast to previous representations such as point- [48, 50, 45], voxel- [30, 49], or mesh-based [65, 17] representations, neural fields do not usually require complex regularization, hand-tuned initialization or optimization control modules as they admit end-to-end optimization and can be queried at arbitrary spatial locations. In the context of view synthesis, Neural Radiance Fields [32] (NeRF) in particular have revolutionized the field by leveraging volumetric rendering, which has proven more robust than prior surface-based rendering approaches [40, 74, 61]. In this work, we employ the robustness and simplicity of neural fields to enable real-time rendering for complex scenes at high quality. More specifically, we use the state-of-the-art radiance field Zip-NeRF [2] to act as a robust prior and source of reliable supervision to train a point-based representation better suited for real-time rendering.
@@
Neural Fields for Real-Time Rendering. NeRF-based models lead to state-of-the-art view synthesis but are typically slow to render so that a variety of works are proposed for speeding up training and inference. While neural fields were initially built on large, compute-heavy multi-layer perceptrons (MLPs) [31, 42, 6, 32], recent works propose the use of voxel representations and interpolation to enable fast training and rendering [12, 77, 55, 35]. Instant NGP [35], for example, demonstrates that a multi-resolution hash grid backbone enables higher quality whilst reducing training time to seconds. However, these works rely on powerful GPUs and often do not achieve real-time rendering for arbitrary scenes. Another line of work aims to represent neural fields as meshes, either as a post-processing step [53, 63] or by direct optimization [73, 18, 5, 66, 54]. These approaches can achieve high frame rates but their quality lags behind volumetric approaches. More recently, another line of work [19, 56, 10, 13] aims to represent a neural field as a set of easily-cacheable assets such as sparse voxel grids, triplanes, and occupancy grids. These methods retain their high quality but often exhibit large storage requirements, are slower to render on smaller devices, and rely on complex custom rendering implementations [10]. In contrast, we optimize lightweight point-based representations that achieve state-of-the-art quality, are easily compressed, and naturally integrate with graphics software following a rasterization pipeline.
@@
Point-Based Representations. First works propose to render point sets as independent geometry samples [15, 16], which can be implemented efficiently in graphics software [57] and highly parallelized on GPU hardware [22, 58]. To eliminate holes when rendering incomplete surfaces, a line of works explores the “splatting” of points with extents larger than a pixel, e.g. with circular or elliptic shapes [67, 76]. The recent work 3D Gaussian Splatting (3DGS) [21] achieves unprecedented quality and fast training and rendering speed by introducing adaptive density control in combination with efficient rasterization kernels. As consequence, 3DGS is used in a variety of applications, including 3D human [79] and avatar reconstruction [9, 51, 33], 3D generation [26, 62, 75, 8], SLAM systems [70, 24, 29], 4D reconstruction [68], and open-set segmentation [60, 52]. Further, works are proposed to address aliasing [78, 71] and point densification [7] in the 3DGS representation. Finally, a recent line of works investigate compression for 3DGS [11, 23, 34, 37, 36] leading to more compact scenes and faster rendering. In this work, we combine a NeRF prior for stable optimization with a point-based 3DGS representation for real-time rendering of complex scenes. Compared to prior works, we enable high-quality view synthesis even for complex real-world captures that might contain lighting and exposure variations. Further, we develop pruning and test-time visibility rendering strategies leading to 10 fewer Gaussian primitives at higher quality compared to 3DGS and with inference times of 900+ FPS.
@@ 3.1 Neural Radiance Fields as a Robust Prior
where are batches of rays sampled from the set of all pixels / rays . To further boost training time and quality, Zip-NeRF [2] uses multisampling and an efficient multi-resolution grid backbone [35]. Due to the state-of-the-art performance, we adopt Zip-NeRF as our radiance field prior.
Robust Optimization on Real-World Data. Real-world captures often contain effects such as lighting and exposure variation or motion blur. Crucial for the success of neural fields on such in-the-wild data [28] is the use of Generative Latent Optimization [3] (GLO) embedding vectors or related techniques. More specifically, a per-image latent vector is optimized along with the neural field that enables explaining away these view-dependent effects
@@ 3.2 Radiance Field-Informed Gaussian Splatting
Gaussian Splatting. In contrast to neural fields, in 3D Gaussian Splatting [21] an explicit point-based scene representation is optimized. More specifically, the scene is represented as points that are associated with a position , opacity , third-degree spherical harmonics (SH) coefficients , 3D scale , and the 3D rotation represented by 4D quaternions . Similar to (1), such a representation can be rendered to the image plane for a camera and a list of correctly-sorted points as
@@
We propose to use the radiance field prior for obtaining a suitable initialization. More specifically, for each pixel / ray from the set of all rays , we first render the median depth from our NeRF model
where are the ray sampling points. We project all pixels / rays into 3D space to obtain our initial point set
where are uniformly randomly-sampled indices for the list of all rays / pixels, indicates the ray origin and the normalized ray direction. We found the median depth estimation to perform better than other common choices such as expected depth by being exact sampling point estimates, and we found setting to million for all scenes to work well in practice. Further, we initialize
and set and to the identity rotation.111We set only as we found to progressively optimize leads to better results [21]. Thus, for each scene we optimize
@@
where is the uniform distribution and we use the default value . Another practical benefit of this approach is that we can train from arbitrary camera lens types due to NeRF’s flexible ray casting, while the 3D Gaussian Splatting gradient formulation assumes a pinhole camera model and it is unclear how this can be efficiently extended to e.g. fisheye or more complex lens types.
@@ 3.3 Ray Contribution-Based Pruning
where indicates the ray contribution for the pixel’s final color prediction in (4) of Gaussian along ray . We find that this formulation leads to improved results compared to concurrent works that investigate similar ideas[23, 11] as we use the exact ray contribution (as opposed to e.g. the opacity) as well as the max operator (instead of e.g. the mean) which is independent of the number of input images, hence more robust to different types of scene coverage.
@@ 3.4 Viewpoint-Based Visibility Filtering
Our pruning technique ensures a compact scene representation with a small overall point count. To scale to larger, more complex scenes such as entire houses or apartments, we introduce a novel viewpoint-based filtering as post-processing step that further speeds up test-time rendering without a quality drop.
@@
Visibility List-Based Rendering. To render an arbitrary viewpoint, we first assign its camera center to the nearest cluster center . Next, we select the respective indicator list . Finally, we perform default rasterization while only considering the points that are marked as active for the respective cluster. This results in a significant FPS increase without any drop in quality.
@@ 4 Experiments
Metrics and Evaluation. We follow common practice and report the view synthesis metrics PSNR, SSIM, and LPIPS. While using techniques such as GLO vectors is essential for high quality on real-world captures (see Sec. 3.1), the evaluation of such models is an open problem such that recent methods [10, 2, 1] train two separate models, one for visualizations, and one (without GLO vectors) purely for the quantitative comparison. In this work, we always train a single model that is robust thanks to the radiance field prior. For evaluation, we simply finetune the trained models on the original image data to match potential color shifts and to ensure a fair comparison. Next to measuring quality, we report the rendering speed in frames per second (FPS) on a RTX 3090 GPU and the number of Gaussians in the scenes (only applicable for point-based methods).
@@ 4.1 Real-Time View Synthesis
Large-Scale Scenes. For the Zip-NeRF dataset [2], we observe a similar trend in Tab. 1(b). Our default and lightweight variant outperform top-performing real-time SMERF and non-real-time Zip-NeRF in SSIM while rendering significantly faster. Notably, our lightweight variant achieves high quality with a mean SSIM of while rendering on average at 748 FPS. In contrast, the state-of-the-art real time method for large scenes, i.e. the large variant of SMERF [10] with submodels, achieves a slightly lower SSIM of with a rendering speed of 221 FPS. Also qualitatively in Fig. 3, we observe that our model achieves high visual appeal with sharper and more stable reconstructions. In contrast to 3DGS [21], we find that our method is more robust on challenging captures as shown in Fig. 4 where 3DGS leads to heavily degraded results on the “Alameda” scene. Note that results still contain floating artifacts, even when equipped with a per-image module that can handle exposure and lighting variations [21, 10]. Our method enables high-quality synthesis even for in-the-wild data.
@@ 4.2 Ablation Study and Limitations
| SSIM | PSNR | LPIPS | # Gaus. (M) | |
@@
Pruning. In Tab. 6(i), we find that our model without pruning leads to similar quantitative results while exhibiting a significantly larger point count. As a result, our pruning technique enables more compact scene representations while maintaining high quality. In Fig. 7 we show that we can match the quality of 3DGS on the “Bicycle” scene, despite having roughly 10 less Gaussians in the scene. Further, we observe a faster increase in SSIM over the first iterations such that our default model can match the 3DGS quality even after only 8k iterations.
| SSIM | PSNR | LPIPS | FPS | |
| Ours | 0.843 | 28.14 | 0.171 | 410 |
| w/o Vis. Fil. | 0.843 | 28.14 | 0.171 | 373 |
| Ours Light | 0.826 | 27.56 | 0.213 | 907 |
| w/o Vis. Fil. | 0.826 | 27.56 | 0.213 | 887 |
| SSIM | PSNR | LPIPS | FPS | |
| Ours | 0.839 | 26.17 | 0.364 | 630 |
| w/o Vis. Fil. | 0.839 | 26.17 | 0.364 | 435 |
| Ours Light | 0.838 | 26.11 | 0.368 | 748 |
| w/o Vis. Fil. | 0.838 | 26.11 | 0.368 | 607 |
@@
Limitations. We optimize a radiance field and a 3DGS representation. While achieving state-of-the-art performance with 900+ FPS, our training time (approx. 2h) is longer compared to single-representation models (0.5h - 1h). Further, despite outperforming prior works on the MipNeRF360 dataset, we observe a small gap to ZipNeRF in PSNR and LPIPS on the large-scale ZipNeRF scenes. We aim to investigate how to achieve faster training and higher quality on large house- and district-level scenes.
@@ 5 Conclusion
We presented RadSplat, a method combining the strengths of radiance fields and Gaussian Splatting for robust real-time rendering of complex scenes with 900 frames per second. We demonstrated that using radiance fields to act as a prior and supervision signal leads to improved results and more stable optimization of point-based 3DGS representations. Our novel pruning technique leads to more compact scenes with a significantly smaller point count, whilst improving quality. Finally, our novel test-time filtering further improves rendering speed without a quality drop. We showed that our method achieves state-of-the-art on common benchmarks while rendering up to 3,000 faster than prior works.
Acknowledgements. We would like to thank Georgios Kopanas, Peter Zhizhin, Peter Hedman, and Jon Barron for fruitful discussions and advice, Cengiz Oztireli for reviewing the draft, and Zhiwen Fan and Kevin Wang for sharing additional baseline results.

### v2

# RadSplat: Radiance Field-Informed Gaussian Splatting for Robust Real-Time Rendering with 900+ FPS

@@ 1 Introduction
While neural field methods have achieved unprecedented view synthesis quality even for challenging real-world captures [2, 37, 33], most approaches are limited by the high compute costs of volumetric rendering. In order to achieve real-time frame rates, recent works reduce network complexity [16, 40], cache intermediate outputs [60, 11], or extract 3D meshes [83, 46, 84, 58]. Nevertheless, all methods trade reduced quality and increased storage costs for faster rendering, and are incapable of maintaining state-of-the-art quality in real-time – the goal of this work (see Fig. 1).
Recently, rasterization-based 3D Gaussian Splatting (3DGS) [26] has emerged as a natural alternative to neural fields. The representation admits real-time frame rates with view synthesis quality rivaling the state-of-the-art in neural fields. 3DGS, however, suffers from a challenging optimization landscape and an unbounded model size. The number of Gaussian primitives is not known a priori, and carefully-tuned merging, splitting, and pruning heuristics are required to achieve satisfactory results. The brittleness of these heuristics become particularly evident in large scenes where phenomena such as exposure variation, motion blur, and moving objects are unavoidable (see Fig. 2). An increasing number of primitives further leads to a potentially-unmanageable memory footprint and reduced rendering speed, strongly limiting model quality for larger scenes.
@@
Our method exhibits state-of-the-art reconstruction quality on both medium and large scenes, with PSNR up to 1.87 dB higher than 3DGS and SSIM exceeding Zip-NeRF, the current state-of-the-art in offline view synthesis (see Fig. 1). At the same time, our method renders up to 907 frames per second, over 3.6 faster than 3DGS [26] and more than 3,000 faster than Zip-NeRF [2].
@@ 2 Related Work
Neural Fields. Since their introduction in the context of 3D reconstruction [36, 47, 5, 78], neural fields have become one of the most promising methods for many 3D vision tasks including 3D/4D reconstruction [84, 30, 49, 44, 48, 50], 3D generative modeling [65, 43, 52, 4, 32, 70], and view synthesis [37, 1, 2]. Key to their success is among others simplicity, state-of-the-art performance, and robust optimization [78, 73, 33]. In contrast to previous representations such as point- [54, 55, 51], voxel- [35, 53], or mesh-based [72, 22] representations, neural fields do not usually require complex regularization, hand-tuned initialization or optimization control modules as they admit end-to-end optimization and can be queried at arbitrary spatial locations. In the context of view synthesis, Neural Radiance Fields [37] (NeRF) in particular have revolutionized the field by leveraging volumetric rendering, which has proven more robust than prior surface-based rendering approaches [45, 82, 67]. In this work, we employ the robustness and simplicity of neural fields to enable real-time rendering for complex scenes at high quality. More specifically, we use the state-of-the-art radiance field Zip-NeRF [2] to act as a robust prior and source of reliable supervision to train a point-based representation better suited for real-time rendering.
@@
Neural Fields for Real-Time Rendering. NeRF-based models lead to state-of-the-art view synthesis but are typically slow to render so that a variety of works are proposed for speeding up training and inference. While neural fields were initially built on large, compute-heavy multi-layer perceptrons (MLPs) [36, 47, 5, 37], recent works propose the use of voxel representations and interpolation to enable fast training and rendering [16, 87, 59, 40]. Instant NGP [40], for example, demonstrates that a multi-resolution hash grid backbone enables higher quality whilst reducing training time to seconds. However, these works rely on powerful GPUs and often do not achieve real-time rendering for arbitrary scenes. Another line of work aims to represent neural fields as meshes, either as a post-processing step [58, 68] or by direct optimization [84, 23, 75, 71, 6, 74, 61]. These approaches can achieve high frame rates but their quality lacks behind volumetric approaches. More recently, another line of work [24, 60, 11, 17] aims to represent a neural field as a set of easily-cacheable assets such as sparse voxel grids, triplanes, and occupancy grids. These methods retain their high quality but often exhibit large storage requirements, are slower to render on smaller devices, and rely on complex custom rendering implementations [11]. In contrast, we optimize lightweight point-based representations that achieve state-of-the-art quality, are easily compressed, and naturally integrate with graphics software following a rasterization pipeline.
@@
Point-Based Representations. First works propose to render point sets as independent geometry samples [20, 21], which can be implemented efficiently in graphics software [62] and highly parallelized on GPU hardware [27, 64]. To eliminate holes when rendering incomplete surfaces, a line of works explores the “splatting” of points with extents larger than a pixel, e.g. with circular or elliptic shapes [76, 86]. The recent work 3D Gaussian Splatting (3DGS) [26] achieves unprecedented quality and fast training and rendering speed by introducing adaptive density control in combination with efficient rasterization kernels. As a consequence, 3DGS is used in a variety of applications, including 3D human [89] and avatar reconstruction [10, 56, 38], 3D generation [31, 69, 85, 8], SLAM systems [80, 29, 34, 63], 4D reconstruction [77, 79], and open-set segmentation [66, 57]. Further, works are proposed to address aliasing [88, 81] and point densification [7, 15] in the 3DGS representation. Finally, a recent line of works investigate compression for 3DGS [12, 28, 39, 42, 41, 13] and the concurrent work [14] analyzes initialization alternatives. In this work, we combine a NeRF prior for stable optimization with a point-based 3DGS representation for real-time rendering of complex scenes. Compared to prior works, we enable high-quality view synthesis even for complex real-world captures that might contain lighting and exposure variations. Further, we develop pruning and test-time visibility rendering strategies leading to 10 fewer Gaussian primitives at higher quality compared to 3DGS and with inference times of 900+ FPS.
@@ 3.1 Neural Radiance Fields as a Robust Prior
where are batches of rays sampled from the set of all pixels / rays . To further boost training time and quality, Zip-NeRF [2] uses multisampling and an efficient multi-resolution grid backbone [40]. Due to the state-of-the-art performance, we adopt Zip-NeRF as our radiance field prior (see supp. mat. for a comparison to an INGP [40] backbone).
Robust Optimization on Real-World Data. Real-world captures often contain effects such as lighting and exposure variation or motion blur. Crucial for the success of neural fields on such in-the-wild data [33] is the use of Generative Latent Optimization [3] (GLO) embedding vectors or related techniques. More specifically, a per-image latent vector is optimized along with the neural field that enables explaining away these image- and view-dependent effects
@@ 3.2 Radiance Field-Informed Gaussian Splatting
Gaussian Splatting. In contrast to neural fields, in 3D Gaussian Splatting [26] an explicit point-based scene representation is optimized. More specifically, the scene is represented as points that are associated with a position , opacity , third-degree spherical harmonics (SH) coefficients , 3D scale , and the 3D rotation represented by 4D quaternions . Similar to (LABEL:eq:render-nerf), such a representation can be rendered to the image plane for a camera and a list of correctly-sorted points as
@@
We propose to use the radiance field prior for obtaining a suitable initialization. More specifically, for each pixel / ray we first define the median depth of our NeRF model as the distance to the first sample along the ray with accumulated transmittance . We unproject all pixels / rays into 3D space to obtain our initial point set
where are uniform randomly-sampled indices for the list of all rays / pixels, indicates the ray origin and the normalized ray direction. We found the median depth estimation to perform better than other common choices such as expected depth by being exact sampling point estimates, and we found setting to million for all scenes to work well in practice (see Sec. 3.2 of supp. mat. for more details). Further, we initialize
and set and to the identity rotation. Thus, for each scene we optimize
@@
where is drawn from the uniform distribution and we use the default value . Another practical benefit of this approach is that we can train from arbitrary camera lens types due to NeRF’s flexible ray casting, while the 3D Gaussian Splatting gradient formulation assumes a pinhole camera model and it is unclear how this can be efficiently extended to e.g. fisheye or more complex lens types.
@@ 3.3 Ray Contribution-Based Pruning
where indicates the ray contribution for the pixel’s final color prediction in (4) of Gaussian along ray . We find that this formulation leads to improved results compared to concurrent works that investigate similar ideas [28, 12] as we use the exact ray contribution (as opposed to e.g. the opacity) as well as the max operator (instead of e.g. the mean) which is independent of the number of input images, hence more robust to different types of scene coverage [19].
@@ 3.4 Viewpoint-Based Visibility Filtering
Our pruning technique ensures a compact scene representation with a small overall point count. To scale to larger, more complex scenes such as entire houses or apartments, inspired by classical occlusion culling [9], we introduce a novel viewpoint-based filtering as post-processing step that further speeds up test-time rendering without a quality drop.
@@
Visibility List-Based Rendering. To render an arbitrary viewpoint, we first assign its camera center to the nearest cluster center . Next, we select the respective indicator list . Finally, we perform default rasterization while only considering the points that are marked as active for the respective cluster. This results in a significant FPS increase up to 45% without any drop in quality.
@@ 4 Experiments
Metrics and Evaluation. We follow common practice and report the view synthesis metrics PSNR, SSIM, and LPIPS. While using techniques such as GLO vectors is essential for high quality on real-world captures (see Sec. 3.1), the evaluation of such models is an open problem such that recent methods [11, 2, 1] train two separate models, one for visualizations, and one (without GLO vectors) purely for the quantitative comparison. In this work, we always train a single model that is robust thanks to the radiance field prior. For evaluation, we simply finetune the trained models on the original training set image data to match potential color shifts and to ensure a fair comparison. Next to measuring quality, we report the rendering speed in frames per second (FPS) on a RTX 3090 GPU and the number of Gaussians in the scenes (only applicable for point-based methods).
@@ 4.1 Real-Time View Synthesis
Large-Scale Scenes. For the Zip-NeRF dataset [2], we observe a similar trend in Tab. 1(b). Our default and lightweight variant outperform top-performing real-time SMERF and non-real-time Zip-NeRF in SSIM while rendering significantly faster. Notably, our lightweight variant achieves high quality with a mean SSIM of while rendering on average at 748 FPS. In contrast, the state-of-the-art real time method for large scenes, i.e. the large variant of SMERF [11] with submodels, achieves a slightly lower SSIM of with a rendering speed of 221 FPS. Also qualitatively in Fig. 4, we observe that our model achieves high visual appeal with sharper and more stable reconstructions. In contrast to 3DGS [26], we find that our method is more robust on challenging captures as shown in Fig. 2 where 3DGS leads to heavily degraded results on the Alameda scene. Note that for 3DGS, results still contain floating artifacts, even when equipped with a per-image module that can handle exposure and lighting variations [26, 11]. In contrast, our method enables high-quality synthesis even for in-the-wild data.
@@ 4.2 Ablation Study and Limitations
| SSIM | PSNR | LPIPS | #G (M) | |
@@
Pruning. In Tab. 6(i), we find that not pruning leads to a small performance drop while exhibiting a significantly larger point count. We hypothesize that a small pruning threshold removes redundancy leading to better generalization. In summary, our pruning technique enables more compact scene representations while maintaining high quality. In Fig. 7 we show that we can match the quality of 3DGS, despite having roughly 10 less Gaussians in the scene. In Fig. 8, we observe a faster increase in SSIM over the first iterations such that we can match the final 3DGS quality after only 8k iterations.
| mip-NeRF 360 dataset | ZipNeRF dataset | |||||||
| SSIM | PSNR | LPIPS | FPS | SSIM | PSNR | LPIPS | FPS | |
| Ours | 0.843 | 28.14 | 0.171 | 410 | 0.839 | 26.17 | 0.364 | 630 |
| w/o Vis. Fil. | 0.843 | 28.14 | 0.171 | 373 | 0.839 | 26.17 | 0.364 | 435 |
| Ours Light | 0.826 | 27.56 | 0.213 | 907 | 0.838 | 26.11 | 0.368 | 748 |
| w/o Vis. Fil. | 0.826 | 27.56 | 0.213 | 887 | 0.838 | 26.11 | 0.368 | 607 |
@@
Limitations. Despite outperforming prior real-time methods, we observe a small gap to ZipNeRF on large-scale scenes which we aim to investigate in the future.
@@ 5 Conclusion
We presented RadSplat, a method combining the strengths of neural fields and point-based representations for robust real-time rendering of complex scenes. Using radiance fields as a prior and supervision signal leads to improved results and more stable optimization of point-based 3DGS representations. Our novel pruning leads to more compact scenes with a significantly smaller scene size, whilst improving quality. Finally, our novel test-time filtering further improves rendering speed without a quality drop. Our method achieves state-of-the-art on common benchmarks while rendering up to 3,000 faster than prior works.